OpenStudy (anonymous):

1. Write the equation of the line which passes through (5, –2) and is parallel to x = 4. (1 point) 2. Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2. (1 point) 3. Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form. (2 point) 4. Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form. (2 point) 5. Using complete sentences, describe one example of a place in your everyday life of parallel lines and one example of perpendicu

OpenStudy (zale101):

use the slope-intercept formula y = mx + b

OpenStudy (precal):

Parallel is same slope, perpendicular is opposite reciprocals. You really should post your questions one at a time.

OpenStudy (anonymous):

Zale101: can you please explain number one? I dont understand at all. I know the answer is 5 but dont know how to find it.

OpenStudy (zale101):

let's say x=4 vertical line that means the line that is parallel line is vertical. x = 4 is a vertical line through 4 on the x-axis, what is parallel to x=4? is x=5

OpenStudy (anonymous):

Okay i get that, so then what is the equation of the line? Sorry i just suck at math.

OpenStudy (zale101):

that's fine we all struggle at it: point slope formula (given point (x1,y1) (y-y1) = m(x-x1)

OpenStudy (zale101):

plug in the numbers and then solve :)

OpenStudy (anonymous):

What about m?

OpenStudy (zale101):

m=slope

OpenStudy (anonymous):

Yeah but what do i put for m? Or how do i solve for it?

OpenStudy (zale101):

in the first problem, you need to graph so you can indicate what is the vertical line that goes through (5,-2), and there's only one line that can do that. x=4 is just a vertical line passing through (4,0). A parallel line to that can also be a vertical line.

OpenStudy (zale101):

There's only one vertical line that goes through (5,-2), and that has to be x=5.

OpenStudy (precal):

problem one using a graph

OpenStudy (precal):